# Advances in Robust Federated Learning: Heterogeneity Considerations

### v1

@@ IV-B2 Functional Encryption
In federated learning, the function $f$ corresponding to the key is often set as Multi-Input Functional Encryption [187], as in HybridAlpha [188]. This setting can effectively prevent inference attacks. In response to the security issues of the TTP, Chotard et al. [189] proposed a decentralized scheme, DMCFE, which allows the system to operate without any TTP, thereby enhancing security. Building on this, 2DMCFE [190] combines the advantages of previous research and can effectively resist mix-and-match attacks.
IV-B3 Secrecy Sharing
Secret sharing (SS) [191, 192] is a technique for splitting and storing secret information, which can effectively prevent the leakage of secrets due to collusion among multiple participants. The core idea is to split the secret into $n$ parts, and at least $t$ slices are needed to fully restore the secret. This scheme is called $(t,n)$-secret sharing.
The most classic in SS is the Shamir scheme [191]. In this scheme, the secret is encoded as a $(t-1)$-dimensional polynomial function $P(x)$, and each participant is assigned the information of a point on $P(x)$. In this way, $t$ participants can restore $P(x)$ from the known information of $t$ points according to the Lagrange interpolation method, thereby restoring the secret. The role of the Shamir scheme in federated learning has been discovered by some researchers, such as the SecAgg [193] and SecAgg+ [194] schemes. The SAFELearn [195] scheme and others extend SS to a multi-server situation. FedShare [196], while applying multi-server SS, also optimizes for utility loss and communication costs.
IV-C Model Tracing
In contrast to the methods previously introduced, the methods discussed in this section primarily serve the function of post-hoc tracing. That is, once an attack has been detected, these methods allow for the identification and verification of the source of the attack. While these methods cannot prevent an attack from occurring in advance, they can obtain evidence after the fact, providing administrators or users of the federated learning system with an advantage at the level of rules or law.
IV-C1 Digital watermarking
@@ IV-C2 Blockchains
Several researchers have proposed numerous federated learning frameworks that incorporate blockchain. Among them, CREAT [203] uses smart contracts on the blockchain to verify the reliability of model information uploaded by clients. Toyoda and Zhang [204] proposed a blockchain federated learning algorithm with an incentive mechanism, which scores each client through a voting scheme to achieve competition among clients. BFLC [142], based on blockchain, implements a decentralized federated learning committee mechanism that can be used to resist malicious attacks. Beyond theoretical research, open-source frameworks like VeryFL [205] are advancing this field from a practical perspective.
@@ V Future Research Directions
With the rapid development of big data and artificial intelligence technologies, people are paying increasing attention to privacy data containing a large amount of potential value. As a distributed machine learning paradigm that protects privacy, FL is gradually becoming an important tool for processing distributed data. In heterogeneous scenarios, FL faces multiple challenges, including data heterogeneity, model heterogeneity, task heterogeneity, communication heterogeneity, device heterogeneity, and privacy protection defects. Despite the existence of the above-mentioned research methods, there are still some challenges that have not been resolved. In the following, we will summarize some potential research directions as follows:
@@
Heterogeneity: (1) Data Distribution [206]: Data heterogeneity is one of the main issues faced by FL in heterogeneous scenarios. Future research can start from the data itself, focusing on preprocessing data across clients and designing more intelligent data selection and filtering mechanisms to improve data quality and model training efficiency. At the same time, we can consider data generalization theory to build a bridge between the source domain and the target domain, thus alleviating the problem of data heterogeneity. (2) Models [78]: Model heterogeneity often leads to differences in model structure, parameter scale, and so on. Future research can focus on model integration and fusion techniques, unifying the semantic information of models to ensure that multiple models with different structures have the same semantic branches, supporting low-loss splicing and fusion of models with different structures. Additionally, for model heterogeneity issues, we can explore optimization techniques such as adaptive training adjustments and model pruning to improve the training efficiency and performance of the model. (3) Tasks [207]: Future research can focus on the analysis and utilization of task relevance, improving the generalization ability of the model by exploring the inherent connections between tasks. Furthermore, we can explore dynamic adjustment and adaptation techniques for model tasks to meet the needs of different tasks and improve the adaptability of the model. (4) Communication [208]: Federated learning relies on the robustness of communication. Studying adaptive network topologies can help mitigate the impact of communication heterogeneity on the central server and improve the overall performance of FL. Additionally, we can focus on the design of fault-tolerant mechanisms for FL to mitigate the negative impact of clients being offline on the overall system. Strategies such as optimizing data transmission formats, reducing communication rounds, and designing efficient communication scheduling strategies are also viable solutions. (5) Devices [56]: Device heterogeneity is mainly reflected in differences in the types of devices participating in FL, computing capabilities, storage capacity, and other aspects. Studying the partitioning and distribution strategies of model parameters and techniques such as model compression and quantization will be feasible solutions to address device heterogeneity issues.
@@
Transfer Learning: Transfer learning can transfer knowledge from one domain to another, thereby accelerating model training in new domains [212]. In federated scenarios, data may be generated in real-time. Therefore, studying how to use FL to achieve cross-domain transfer learning has important practical significance for improving the adaptability and performance of FL.
@@
VI Conclusion
In this paper, we provide insights into key concepts in the field of heterogeneous federated learning and the multiple challenges it faces. In particular, we analyze in detail the heterogeneity of data, models, tasks, communications, and devices, which pose numerous algorithmic challenges in federated learning systems. To address these challenges, in this paper, we systematically categorize existing research approaches by classifying them into data-level, model-level, and structure-level according to the processing hierarchy, each of which corresponds to a specific problem and solution strategy. Our categorization approach provides a clear perspective for understanding the efficacy of various types of algorithms in coping with heterogeneous environments. With this hierarchical strategy, researchers and practitioners can more efficiently select methods that fit their specific needs. Finally, this paper looks at future research directions in heterogeneous federated learning, which will undoubtedly continue to be a dynamic and challenging research direction in the field of machine learning as technology advances and application needs increase. We look forward to more innovative research results in this field in the future to promote the development of technology and popularization of applications.
Acknowledgment
The research is supported by the National Key Research and Development Program of China (2023YFB2703700), the National Natural Science Foundation of China (62176269), the Guangzhou Science and Technology Program (2023A04J0314), the National Natural Science Foundation of China and Guangdong Provincial Joint Fund (No. U1911202).

### v2

@@ IV-C2 Functional Encryption
In federated learning, the function $f$ corresponding to the key is often set as Multi-Input Functional Encryption [192], as in HybridAlpha [73]. This setting can effectively prevent inference attacks. In response to the security issues of the TTP, Chotard et al. [193] proposed a decentralized scheme, DMCFE, which allows the system to operate without any TTP, thereby enhancing security.
IV-C3 Secrecy Sharing
Secret sharing (SS) [194] is a technique for splitting and storing secret information, which can effectively prevent the leakage of secrets due to collusion among multiple participants. The core idea is to split the secret into $n$ parts, and at least $t$ slices are needed to fully restore the secret. This scheme is called $(t,n)$-secret sharing. The most classic in SS is the Shamir scheme, which is built upon the Lagrange interpolation method. The role of the Shamir scheme in federated learning has been discovered by some researchers, such as the SecAgg [195] and SecAgg+ [196] schemes.
IV-C4 SMC with Heterogeneiry
Although SMC can provide a high level of security to FL, it is still negatively affected by the heterogeneity of FL nodes. The biggest challenge lies in how different clients can generate the same key in heterogeneous networks.
Trustworthy Third Party (TTP) [197]. Clearly, clients can jointly trust an external node as a TTP, which generates and securely distributes the key to each client. This process is known as key distribution. Introducing a TTP to overcome heterogeneity is a simple method that requires minimal computational overhead. However, it relies on complete trust in the TTP. If the TTP itself is malicious, the protection provided by SMC will be compromised.
Key Agreement Protocols [198]. In this protocol, the key is not centrally generated; instead, each participating node exchanges the necessary information to independently compute the same session key. Similar to common encryption schemes, key agreement protocols are implemented through mathematically challenging problems, ensuring that adversaries, even upon intercepting the communication, cannot infer the session key. However, as a trade-off for security, this method incurs a significant increase in computational overhead.
Both of the aforementioned solutions have their respective advantages and disadvantages, necessitating that managers select based on the specific circumstances of the system. Generally, TTP emphasizes efficiency, whereas the key agreement protocol places a greater emphasis on security.
IV-D Model Tracing
In contrast to the methods previously introduced, the methods discussed in this section primarily serve the function of post-hoc tracing. That is, once an attack has been detected, these methods allow for the identification and verification of the source of the attack. While these methods cannot prevent an attack from occurring in advance, they can obtain evidence after the fact, providing administrators or users of the federated learning system with an advantage at the level of rules or law. Most model tracing methods are transparent to FL systems, functioning as plugins that seamlessly integrate into the training process while avoiding the limitations posed by the heterogeneous environment.
IV-D1 Digital watermarking
@@ IV-D2 Blockchains
Several researchers have proposed numerous federated learning frameworks that incorporate blockchain. Among them, CREAT [77] uses smart contracts on the blockchain to verify the reliability of model information uploaded by clients. BFLC [78], based on blockchain, implements a decentralized federated learning committee mechanism that can be used to resist malicious attacks. Beyond theoretical research, open-source frameworks like VeryFL [79] are advancing this field from a practical perspective.
@@ V Future Research Directions
With the rapid development of big data and artificial intelligence technologies, FL, as a privacy-preserving distributed machine learning paradigm, is becoming an important tool for processing distributed data. In heterogeneous scenarios, FL faces multiple heterogeneous challenges. Despite the existence of the above-mentioned research methods, there are still some challenges that have not been resolved. Previous solutions usually suffer from the lack of granularity of heterogeneous alignment granularity, the need for additional computational and communication costs, the relaxation of privacy-preserving assumptions. Furthermore, there is a lack of consideration of special settings such as heterogeneous models, heterogeneous devices, fairness, and so on. Therefore, in the following, we will summarize some potential research directions as follows:
@@
Heterogeneity: (1) Data Distribution [202]: Data heterogeneity is one of the main issues faced by FL in heterogeneous scenarios. Future research can start from the data itself, focusing on preprocessing data across clients and designing more intelligent data selection and filtering mechanisms to improve data quality and model training efficiency. At the same time, we can consider data generalization theory to build a bridge between the source domain and the target domain, thus alleviating the problem of data heterogeneity. (2) Models [49]: Model heterogeneity often leads to differences in model structure, parameter scale, and so on. Future research can focus on model integration and fusion techniques, unifying the semantic information of models to ensure that multiple models with different structures have the same semantic branches, supporting low-loss splicing and fusion of models with different structures. Additionally, for model heterogeneity issues, we can explore optimization techniques such as adaptive training adjustments and model pruning to improve the training efficiency and performance of the model. (3) Tasks [203]: Future research can focus on the analysis and utilization of task relevance, improving the generalization ability of the model by exploring the inherent connections between tasks. Furthermore, we can explore dynamic adjustment and adaptation techniques for model tasks to meet the needs of different tasks and improve the adaptability of the model, even to be able to handle federated requirements with completely different tasks. (4) Communication [204]: FL relies on the robustness of communication. Studying adaptive network topologies can help mitigate the impact of communication heterogeneity on the central server and improve the overall performance of FL. Additionally, we can focus on the design of fault-tolerant mechanisms for FL to mitigate the negative impact of clients being offline on the overall system. Strategies such as optimizing data transmission formats, reducing communication rounds, and designing efficient communication scheduling strategies are also viable solutions. (5) Devices [98]: Device heterogeneity is mainly reflected in differences in the types of devices participating in FL, computing capabilities, storage capacity, and other aspects. Studying the partitioning and distribution strategies of model parameters and techniques such as model compression and quantization will be feasible solutions to address device heterogeneity issues.
@@
Foundation Models: Foundation models have demonstrated their potential in a variety of scenarios. However, there are potential privacy concerns due to the large amount of data that needs to be introduced during the training process. FL as a privacy-preserving approach, it would be meaningful work to think about the combination with the foundation model. This combination can be considered from a variety of perspectives, for instance, thinking about the federated data component for the foundation models [208].
Plug and Play Edge Participants: with the development of mobile technology and IoT devices, federated learning participants may not be online all the time due to power and network limitations. The flexibility of the devices makes the joining and dropping of clients frequent. Therefore, it is necessary to investigate federated learning algorithms with flexibility and scalability.